A 25 mL sample of H3PO4 requires 31.15 mL of 0.242 N KOH for its titration. What is the normality of this H3PO4 (aq) if it is always to be used in reaction?
1
Expert's answer
2020-12-01T09:18:31-0500

"C_n(H_3PO_4) *V(H_3PO_4) ="

"= C_n(KOH)*V(KOH)"

"C_n(H_3PO_4) = C_n(KOH)*\frac{V(KOH)}{V(H_3PO_4)} ="

"= 0.242\ N * \frac{31.15\ ml}{25\ ml} ="

"= 0.301532\ N"


Answer: "C_n(H_3PO_4) = 0.301532\ N"